OpenStudy (anonymous):

what is the oxidation number of Cl in ClO3-

OpenStudy (zpupster):

Cl = ?? O has oxidation number -2 and you have 3 so 3(-2) = -6 but you see the ClO3 has an oxidation of -1 s0 -6 + ( the oxidation of Cl) = -1

OpenStudy (cuanchi):

No, sorry, that is not correct, please review the calculations, -6 - (-1) = ? the oxidation of Cl it is not -1 in ClO3-

OpenStudy (zpupster):

no it is +5 see -6+ 5 -=-1

OpenStudy (zpupster):

-6 + (what is the oxidation of Cl) = -1